# Investigating Material Interface Diffusion Phenomena through Graph Neural Networks in Applied Materials

### v2

@@ 1 INTRODUCTION
Materials science is a pivotal discipline driving modern technological advancements, focusing on the composition, structure, properties, and relationships of materials. Within materials engineering, interface diffusion profoundly influences mechanical properties, conductivity, and thermal stability, impacting various practical applications.1, 2 However, traditional research methods, relying on experimental techniques and numerical simulations, often struggle with complex multiphase interface problems due to their time-consuming and labor-intensive nature.3 Despite significant advancements in materials science and engineering, many unresolved challenges persist in the study of interfacial diffusion. The complexity of interfacial phenomena, driven by diverse material properties and environmental conditions, continues to elude comprehensive understanding. Moreover, the heterogeneous nature of materials and the multi-scale interactions involved further complicate accurate prediction and control of diffusion processes at interfaces.
@@
To address these challenges, it is essential to identify specific unresolved issues related to material interface diffusion in various material systems, as detailed in Table. 1. These issues are critical in the context of materials science and engineering, especially in applications involving semiconductors, electronic devices, solid-state batteries, and thin-film technologies.4, 5, 6, 7 For instance, the diffusion at the interface between silicon (Si) and silicon dioxide (SiO2) significantly affects the electrical properties in semiconductor manufacturing, posing challenges in enhancing device performance.8 Similarly, the electromigration phenomenon due to copper (Cu) interconnect diffusion in silicon (Si) impacts circuit stability and resistance, requiring advanced strategies to mitigate these effects.9 In the case of germanium-doped silicon (Ge-Si), understanding the interface diffusion is essential for managing crystal defects and optimizing electronic properties.10 Additionally, multilayer nitride films on silicon need careful examination of diffusion impacts to maintain film stress and structural stability.11 High-power electronic devices benefit from controlled silicon and silicon carbide (SiC) interface diffusion,12 while the dielectric properties and adhesion of aluminum oxide (Al2O3) thin films depend on minimizing interface diffusion with silicon.13 Furthermore, the performance of high-frequency devices can be compromised by the diffusion at the silicon and gallium nitride (GaN) interface.14 The contact resistance and reliability issues arising from gold (Au) diffusion in silicon,15 along with the electrical properties of nickel silicide (NiSi) affected by nickel (Ni) diffusion, highlight the need for precise interface management.11 Finally, optimizing the performance of optoelectronic devices involves addressing silicon and indium phosphide (InP) interface diffusion challenges.16 These examples underscore the importance of advanced research and innovative solutions in managing material interface diffusion to enhance the performance and reliability of various technological applications.
@@
In recent years, the rapid advancement of artificial intelligence, particularly in the realm of machine learning, offers promising avenues for materials science. Graph Neural Networks (GNNs), adept at handling non-euclidean data such as graph-structured data, have emerged as a novel approach with distinct advantages.17 GNNs not only adeptly capture complex relationships within material microstructures but also efficiently predict and optimize material properties.18 In this context, this paper investigates the application of GNN technology to study diffusion phenomena at material interfaces. Through the development of an efficient GNN model, the aim is to achieve more accurate predictions within shorter timeframes, thereby furnishing novel insights and methodologies for materials science research.19, 20
@@
2 COMPUTATIONAL METHODS AND MODELS
@@ 2.1 Fundamentals and advantages of graph neural networks
GNNs represent a sophisticated class of neural network models specifically engineered to operate on graph-structured data. These networks offer a robust framework for encoding intricate relationships between entities within high-dimensional data matrices.21 The foundational strength of GNNs lies in their adeptness at capturing and understanding complex relationships inherent in graph-structured data.
@@
The ability of GNNs to model and learn from high-dimensional structured data, coupled with their computational efficiency, positions them as versatile tools applicable across various domains. Notably, GNNs find extensive utility in applications such as molecular structure prediction and recommender systems, where their capacity to grasp intricate relationships within data proves invaluable.21 For instance, in molecular structure prediction, GNNs can model the interactions between atoms and predict molecular properties with high accuracy. In recommender systems, GNNs can capture user-item interactions to provide personalized recommendations.
@@ 2.2 Case studies of graph neural networks in material property prediction
GNNs demonstrate remarkable proficiency in discerning intricate patterns within molecular structures and accurately forecasting properties such as energy, stability, and reactivity. This surpasses the capabilities of traditional data-driven algorithms, heralding a new era in materials prediction. For instance, recent advancements in materials science have witnessed the emergence of GNNs as formidable instruments for predicting material properties based on atomic structure. These neural network models offer a paradigm shift by representing molecules as graphs, with atoms serving as nodes and bonds as edges, thereby enabling a nuanced description of materials’ microscopic constituents and structural attributes.22, 23
@@
To better understand the advantages of GNNS, it is essential to compare them with traditional simulation methods. Traditional methods, such as molecular dynamics (MD) and density functional theory (DFT), often require extensive computational resources and time, particularly for complex systems with large numbers of atoms and intricate interactions.24 These methods also depend heavily on accurate potential energy functions and may face challenges in capturing long-time scale phenomena due to their inherent limitations.25 In contrast, machine learning algorithms can leverage large datasets to learn and predict complex diffusion behaviors with high accuracy and efficiency. By training on existing experimental and simulation data, machine learning models can generalize to new conditions and materials without the need for exhaustive recalculations. This ability to generalize makes machine learning particularly powerful for predicting diffusion properties across a wide range of materials and interface conditions.26 Additionally, machine learning models can incorporate a broader set of features, including atomic environments, local chemistry, and external conditions, to provide more comprehensive predictions. These models can also be updated and improved continuously as new data becomes available, enhancing their predictive power over time. Moreover, machine learning techniques can uncover hidden patterns and relationships in diffusion processes that may not be apparent through traditional methods, offering novel insights and guiding experimental design more effectively. Overall, the integration of machine learning algorithms into material interface diffusion studies presents a transformative approach, enabling faster, more accurate, and more insightful analysis compared to conventional simulation techniques.
@@
Beyond diffusion studies, the application spectrum of GNNs in materials science extends beyond mere prediction; they play pivotal roles in catalyst design by modeling catalyst atomic structures and interactions with reactant molecules, thus predicting catalytic activity and selectivity. Similarly, in battery material prediction and design, GNNs leverage atomic structure and crystallographic information to predict crucial properties such as capacity, cycle stability, and rate capability, thereby streamlining research efforts in this domain. Furthermore, recent studies underscore the superiority of GNNs over traditional machine learning models in predicting material properties such as bandgaps in semiconductor crystals, thereby offering valuable insights into structure-property relationships and accelerating materials discovery and design endeavors.27
@@ 2.3 Data acquisition and preprocessing
Data Collection: Experimental data are collected from literature, databases, and experimental reports detailing the diffusion coefficients, concentration gradients, and other relevant parameters of different materials. Additionally, we utilize molecular dynamics (MD) simulations and finite element analysis (FEA) to generate supplementary data, providing detailed atomic-level insights into the diffusion processes.28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51 This comprehensive collection ensures a robust dataset for subsequent analysis.
@@
Feature Extraction: Key features that influence interface diffusion, such as atomic radius, electronegativity, lattice parameters, and bonding characteristics, are extracted. To determine the importance of various descriptors in predicting material interface diffusion, we employed Principal Component Analysis (PCA). PCA is a dimensionality reduction technique that transforms the original set of features into a new set of uncorrelated components, capturing the most variance in the data. In our analysis, we applied PCA to a dataset containing descriptors related to material properties and diffusion parameters. By examining the loadings of the original features on these principal components, we assessed the contribution of each descriptor to the overall variance in the diffusion phenomena. The percentage values shown in Figure 1 represent the relative importance of each descriptor, derived from its contribution to the most significant principal components.
This approach allowed us to objectively identify which descriptors were most influential in predicting diffusion behavior, providing a clear understanding of the factors driving the observed phenomena. The use of PCA ensured that the most critical features were highlighted, facilitating a more accurate and interpretable model.
@@ 2.4 Feature engineering and input data preparation
To predict material interface diffusion efficiently using machine learning, we need to handle the identified influential factors appropriately as input features for our model. Key features such as atomic radius and lattice constant are included as input features, using normalized or standardized values as necessary. The vacancy concentration of materials under specific conditions, derived from experimental data or simulations, is also included. Types and concentrations of interstitial atoms are considered as input features. The temperature at which the material is studied is included, normalized or standardized as necessary. Chemical composition is represented using one-hot encoding or embedding vectors, and the types and concentrations of impurities are included as input features. Grain boundary area or grain size and phase boundary area or phase boundary density are also included. The type of crystal structure is encoded as a numerical feature (embedding vectors). Values of mechanical stress experienced by the material and deformation levels or strain are included as input features. The strength and direction of any applied electric field and magnetic field are also considered as input features.50 These features ensure a comprehensive representation of the factors influencing diffusion.
@@ 2.6 Data splitting
The dataset is split into training, validation, and test sets, typically in a ratio of 70:15:15. This ensures that each set is representative of the overall dataset to avoid bias.
3 RESULTS AND DISCUSSION
@@ 3.2 Model validation
To validate the accuracy and reliability of our GNN models, we conducted a series of validation experiments using experimental data from previous studies. These experiments were designed to test our model’s predictions against established datasets that detailed diffusion behavior at various time points across different material interfaces. By comparing the predicted diffusion coefficients, diffusion rates, and concentration profiles generated by our GNN model with actual experimental observations, we assessed the model’s precision. The comparison revealed that our model consistently produced results closely matching the observed data, demonstrating its capability to accurately predict interface diffusion phenomena. This validation process not only confirmed the robustness of our approach but also highlighted its potential for broader applications in material science.
@@ 3.3 Model outputs
To derive the solutions listed in Table 4, we systematically adjusted key input parameters in our GNN models and analyzed their effects on diffusion behavior at various material interfaces. One critical parameter we manipulated was temperature. By increasing the temperature in our simulations, the model predicted a corresponding increase in the diffusion rate of silicon atoms into the SiO2 layer. This prediction highlighted the significant influence of temperature on interface diffusion.
Based on these model predictions, we provided targeted recommendations for optimizing temperature settings to achieve specific diffusion characteristics. Similar adjustments were made to other parameters, such as pressure, chemical composition, and layer thickness, allowing us to observe their impacts and derive practical solutions tailored to meet specific material performance requirements. These solutions were directly informed by the predictive capabilities of our GNN models, ensuring they are both actionable and scientifically grounded.
This approach enabled us to extract insights applicable to real-world material design and optimization, demonstrating the practical utility of our modeling framework.
@@ 4 CONCLUSIONS
Therefore, we propose a novel application of Graph Neural Networks (GNNs) to model material interface diffusion, marking a significant advancement over traditional theoretical models. Traditional models largely depend on predefined assumptions and theoretical frameworks, which often limit their predictive accuracy and adaptability to complex systems 35. In contrast, our approach utilizes a data-driven methodology that allows the GNN to learn and predict diffusion behavior directly from experimental or simulated data, without being constrained by prior assumptions about the system. This flexibility enables our model to achieve superior performance in predicting key diffusion characteristics, such as diffusion coefficients, diffusion rates, and concentration profiles, across a range of material interfaces. Furthermore, our GNN-based model provides a more detailed and accurate mapping of potential diffusion pathways, a task that is challenging for conventional methods due to their inherent limitations.
@@
Figure 2: Workflow of this study, illustrating the interfacing materials of two phases represented as a matrix and input into the GNN model for predicting interface diffusion phenomena. In this figure, (a) shows the actual crystal structure at the contact interface, using Au and Cu as examples. (b) Abstracts this structure into a matrix form for computational modeling. (c) Demonstrates the matrix as input into the model for training purposes. (d) Displays the predicted outcomes derived from the trained model. In the diffusion process, Au and Cu atoms vacate their original lattice positions to facilitate diffusion.
@@
SSEs and cathodes Strategies to reduce interfacial resistance. 4
Si and SiO2 Investigating the impact of SiO2-Si interface diffusion on electrical properties. 8
Cu and Si Examining the electromigration phenomenon due to Cu interconnect diffusion in Si. 9
Ge-Si Assessing how Ge-Si interface diffusion affects crystal defects. 10
Si and ML. Ni Analyzing the impact of nitride diffusion on film stress and structural stability. 11
Si and Ni Optimizing the formation and electrical properties of nickel silicide (NiSi). 11
Si and SiC Investigating Si-SiC interface diffusion to enhance performance in high-power devices. 12
Si and Al2O3 Enhancing dielectric properties and adhesion of Al2O3 thin films. 13
Si and GaN Examining the impact of interface diffusion on high-frequency devices. 14
Si and Au Addressing increased contact resistance issues and reliability due to Au diffusion. 15
Si and InP Investigating the impact of Si-InP interface diffusion on optoelectronic devices. 16
p- and n-type SCs Exploring methods to minimize ID affecting electrical and thermal conductivities. 51
@@
This work was supported by the Science and Technology Development Fund, Macao SAR (File No. 0090/2021/A2), University of Macau (MYRG-GRG2024-00158-IAPME), and the Guangdong-Hong Kong-Macao Joint Laboratory for Neutron Scattering Science and Technology (Grant No. 2019B121205003).

### v3

@@ Abstract
Caption of Graphical Abstract: By analyzing experimental and simulated data, GNNs accurately forecast diffusion coefficients, rates, and pathways, enhancing material design and performance.
@@ 1 Introduction
Materials science is a pivotal discipline driving modern technological advancements, focusing on the composition, structure, properties, and relationships of materials. Within materials engineering, interface diffusion profoundly influences mechanical properties, conductivity, and thermal stability, impacting various practical applications.gupta1998interface, goodenough2010challenges However, traditional research methods, relying on experimental techniques and numerical simulations, often struggle with complex multiphase interface problems due to their time-consuming and labor-intensive nature.ward1952diffusion Despite significant advancements in materials science and engineering, many unresolved challenges persist in the study of interfacial diffusion. The complexity of interfacial phenomena, driven by diverse material properties and environmental conditions, continues to elude comprehensive understanding. Moreover, the heterogeneous nature of materials and the multi-scale interactions involved further complicate accurate prediction and control of diffusion processes at interfaces.
@@
To address these challenges, it is essential to identify specific unresolved issues related to material interface diffusion in various material systems, as detailed in Table. 1. These issues are critical in the context of materials science and engineering, especially in applications involving semiconductors, electronic devices, solid-state batteries, and thin-film technologies.ZHOU2023101248, https://doi.org/10.1002/ejic.202300382, XIAO2024151111, WANG202479 For instance, the diffusion at the interface between silicon (Si) and silicon dioxide (SiO2) significantly affects the electrical properties in semiconductor manufacturing, posing challenges in enhancing device performance.fukatsu2004effect Similarly, the electromigration phenomenon due to copper (Cu) interconnect diffusion in silicon (Si) impacts circuit stability and resistance, requiring advanced strategies to mitigate these effects.takeyama1996properties In the case of germanium-doped silicon (Ge-Si), understanding the interface diffusion is essential for managing crystal defects and optimizing electronic properties.silvestri2006diffusion Additionally, multilayer nitride films on silicon need careful examination of diffusion impacts to maintain film stress and structural stability.chang1982diffusion High-power electronic devices benefit from controlled silicon and silicon carbide (SiC) interface diffusion,komninou1999gold while the dielectric properties and adhesion of aluminum oxide (Al2O3) thin films depend on minimizing interface diffusion with silicon.werner2011electronic Furthermore, the performance of high-frequency devices can be compromised by the diffusion at the silicon and gallium nitride (GaN) interface.lin2000improved The contact resistance and reliability issues arising from gold (Au) diffusion in silicon,slezak2000surface along with the electrical properties of nickel silicide (NiSi) affected by nickel (Ni) diffusion, highlight the need for precise interface management.chang1982diffusion Finally, optimizing the performance of optoelectronic devices involves addressing silicon and indium phosphide (InP) interface diffusion challenges.shapira1984interdiffusion These examples underscore the importance of advanced research and innovative solutions in managing material interface diffusion to enhance the performance and reliability of various technological applications.
@@
In recent years, the rapid advancement of artificial intelligence, particularly in the realm of machine learning, offers promising avenues for materials science. Graph Neural Networks (GNNs), adept at handling non-euclidean data such as graph-structured data, have emerged as a novel approach with distinct advantages.gori2005new GNNs not only adeptly capture complex relationships within material microstructures but also efficiently predict and optimize material properties.bruna2013spectral In this context, this paper investigates the application of GNN technology to study diffusion phenomena at material interfaces. Through the development of an efficient GNN model, the aim is to achieve more accurate predictions within shorter timeframes, thereby furnishing novel insights and methodologies for materials science research.wang2023capacity, verma2018graph
@@
2 Graph neural networks and applications
@@ 2.1 Fundamentals and advantages of graph neural networks
GNNs represent a sophisticated class of neural network models specifically engineered to operate on graph-structured data. These networks offer a robust framework for encoding intricate relationships between entities within high-dimensional data matrices.griggs2020unified The foundational strength of GNNs lies in their adeptness at capturing and understanding complex relationships inherent in graph-structured data.
@@
The ability of GNNs to model and learn from high-dimensional structured data, coupled with their computational efficiency, positions them as versatile tools applicable across various domains. Notably, GNNs find extensive utility in applications such as molecular structure prediction and recommender systems, where their capacity to grasp intricate relationships within data proves invaluable.griggs2020unified For instance, in molecular structure prediction, GNNs can model the interactions between atoms and predict molecular properties with high accuracy. In recommender systems, GNNs can capture user-item interactions to provide personalized recommendations.
@@ 2.2 Case studies of graph neural networks in material property prediction
GNNs demonstrate remarkable proficiency in discerning intricate patterns within molecular structures and accurately forecasting properties such as energy, stability, and reactivity. This surpasses the capabilities of traditional data-driven algorithms, heralding a new era in materials prediction. For instance, recent advancements in materials science have witnessed the emergence of GNNs as formidable instruments for predicting material properties based on atomic structure. These neural network models offer a paradigm shift by representing molecules as graphs, with atoms serving as nodes and bonds as edges, thereby enabling a nuanced description of materials’ microscopic constituents and structural attributes.chen2023md, sathana2022prediction
@@
To better understand the advantages of GNNS, it is essential to compare them with traditional simulation methods. Traditional methods, such as molecular dynamics (MD) and density functional theory (DFT), often require extensive computational resources and time, particularly for complex systems with large numbers of atoms and intricate interactions.liu2004calculation These methods also depend heavily on accurate potential energy functions and may face challenges in capturing long-time scale phenomena due to their inherent limitations.patel2019computing In contrast, machine learning algorithms can leverage large datasets to learn and predict complex diffusion behaviors with high accuracy and efficiency. By training on existing experimental and simulation data, machine learning models can generalize to new conditions and materials without the need for exhaustive recalculations. This ability to generalize makes machine learning particularly powerful for predicting diffusion properties across a wide range of materials and interface conditions.allers2020machine Additionally, machine learning models can incorporate a broader set of features, including atomic environments, local chemistry, and external conditions, to provide more comprehensive predictions. These models can also be updated and improved continuously as new data becomes available, enhancing their predictive power over time. Moreover, machine learning techniques can uncover hidden patterns and relationships in diffusion processes that may not be apparent through traditional methods, offering novel insights and guiding experimental design more effectively. Overall, the integration of machine learning algorithms into material interface diffusion studies presents a transformative approach, enabling faster, more accurate, and more insightful analysis compared to conventional simulation techniques.
@@
Beyond diffusion studies, the application spectrum of GNNs in materials science extends beyond mere prediction; they play pivotal roles in catalyst design by modeling catalyst atomic structures and interactions with reactant molecules, thus predicting catalytic activity and selectivity. Similarly, in battery material prediction and design, GNNs leverage atomic structure and crystallographic information to predict crucial properties such as capacity, cycle stability, and rate capability, thereby streamlining research efforts in this domain. Furthermore, recent studies underscore the superiority of GNNs over traditional machine learning models in predicting material properties such as bandgaps in semiconductor crystals, thereby offering valuable insights into structure-property relationships and accelerating materials discovery and design endeavors.zhang2023graph
3 Research methods
@@ 3.1 Data acquisition and preprocessing
Data Collection: Experimental data are collected from literature, databases, and experimental reports detailing the diffusion coefficients, concentration gradients, and other relevant parameters of different materials. Additionally, we utilize molecular dynamics (MD) simulations and finite element analysis (FEA) to generate supplementary data, providing detailed atomic-level insights into the diffusion processes.jiangwei2002microstructure, svoboda2011diffusion, yang2020comprehensive, janssen1973diffusion, chambers1987silicide, peng2005study, divinski2004ag, erdelyi2002interface, burger1987structure, li2014molecular, ogawa1995generalized, shen2001electronic, lee1997prediction, lee2014two, whitlow1991diffusion, adalsteinsson2003transport, tavoosi2017kirkendall, luthra1991mixed, trumble1991thermodynamics, watanabe1994electron, langer1975theory, griscom1985diffusion This comprehensive collection ensures a robust dataset for subsequent analysis.
@@
Feature Extraction: Key features that influence interface diffusion, such as atomic radius, electronegativity, lattice parameters, and bonding characteristics, are extracted.
To determine the importance of various descriptors in predicting material interface diffusion, we employed Principal Component Analysis (PCA). PCA is a dimensionality reduction technique that transforms the original set of features into a new set of uncorrelated components (principal components), which capture the most variance in the data. In our analysis, we applied PCA to the dataset containing the various descriptors related to material properties and diffusion parameters. By examining the loadings of the original features on these principal components, we were able to assess the contribution of each descriptor to the overall variance in the diffusion phenomena. The percentage values shown in Figure 1 represent the relative importance of each descriptor, as derived from its contribution to the most significant principal components. This approach allowed us to objectively identify which descriptors were most influential in predicting the diffusion behavior, providing a clear understanding of the factors driving the observed phenomena. The use of PCA thus ensured that the most critical features were highlighted, facilitating a more accurate and interpretable model.
@@ 3.2 Feature engineering and input data preparation
To predict material interface diffusion efficiently using machine learning, we need to handle the identified influential factors appropriately as input features for our model. Key features such as atomic radius and lattice constant are included as input features, using normalized or standardized values as necessary. The vacancy concentration of materials under specific conditions, derived from experimental data or simulations, is also included. Types and concentrations of interstitial atoms are considered as input features. The temperature at which the material is studied is included, normalized or standardized as necessary. Chemical composition is represented using one-hot encoding or embedding vectors, and the types and concentrations of impurities are included as input features. Grain boundary area or grain size and phase boundary area or phase boundary density are also included. The type of crystal structure is encoded as a numerical feature (embedding vectors). Values of mechanical stress experienced by the material and deformation levels or strain are included as input features. The strength and direction of any applied electric field and magnetic field are also considered as input features.mackiewicz1993principal These features ensure a comprehensive representation of the factors influencing diffusion.
@@ 3.4 Data splitting
The dataset is split into training, validation, and test sets, typically in a ratio of 70:15:15. This ensures that each set is representative of the overall dataset to avoid bias. Next, we will discuss the model training precess in detail.
4 Experimental section
@@ 4.2 Model validation
Therefore, we conducted a series of validation experiments using experimental data from previous studies. Specifically, we tested our model’s predictions against established datasets that detailed the diffusion behavior at various time points across different material interfaces. By comparing the predicted diffusion coefficients, diffusion rates, and concentration profiles generated by our GNN model with the actual experimental observations, we were able to assess the model’s precision.
The comparison revealed that our model consistently produced results that closely matched the observed data, demonstrating its capability to accurately predict interface diffusion phenomena. This validation process not only confirmed the robustness of our approach but also highlighted its potential for broader applications in material science.
@@ 5 Results and discussion
To derive the solutions listed in Table 4, we systematically adjusted key input parameters in our GNN models and analyzed their effects on the diffusion behavior at various material interfaces. For example, one critical parameter we manipulated was temperature. By increasing the temperature in our simulations, the model predicted a corresponding increase in the diffusion rate of silicon atoms into the SiO2 layer. This prediction highlighted the temperature’s significant influence on interface diffusion. Based on these model predictions, we provided targeted recommendations for optimizing temperature settings to achieve specific diffusion characteristics. Similar adjustments were made to other parameters, such as pressure, chemical composition, and layer thickness, allowing us to observe their impacts and derive practical solutions tailored to meet specific material performance requirements. These solutions were directly informed by the predictive capabilities of our GNN models, ensuring that they are both actionable and scientifically grounded. This approach enabled us to extract insights that are applicable to real-world material design and optimization, demonstrating the practical utility of our modeling framework.
@@ 6 Conclusion
Therefore, we proposed a novel application of Graph Neural Networks (GNNs) to model material interface diffusion, offering a significant advancement over traditional theoretical models. Previous studies predominantly relied on predefined assumptions and theoretical frameworks to describe diffusion processes, which often limited their predictive accuracy and adaptability to complex systems erdelyi2002interface. In contrast, our approach utilized a data-driven methodology that allowed the GNN to learn and predict diffusion behavior directly from experimental or simulated data, without being constrained by prior assumptions about the system.
This flexibility enabled our model to achieve superior performance in predicting key diffusion characteristics, such as diffusion coefficients, diffusion rates, and concentration profiles, across a range of material interfaces. Furthermore, our GNN-based model provided a more detailed and accurate mapping of potential diffusion pathways, a task that was challenging for conventional methods due to their inherent limitations.
@@
Figure 2: Workflow of this study, illustrating the interfacing materials of two phases represented as a matrix and input into the GNN model for predicting interface diffusion phenomena. In this figure, (a) shows the actual crystal structure at the contact interface, using Au and Cu as examples. (b) Abstracts this structure into a matrix form for computational modeling. (c) Demonstrates the matrix as input into the model for training purposes. (d) Displays the predicted outcomes derived from the trained model, Where Au, Cu atoms left their original lattice positions to complete diffusion.
@@
SSEs and cathodes Strategies to reduce interfacial resistance. ZHOU2023101248
Si and SiO2 Investigating the impact of SiO2-Si interface diffusion on electrical properties. fukatsu2004effect
Cu and Si Examining the electromigration phenomenon due to Cu interconnect diffusion in Si. takeyama1996properties
Ge-Si Assessing how Ge-Si interface diffusion affects crystal defects. silvestri2006diffusion
Si and ML. Ni Analyzing the impact of nitride diffusion on film stress and structural stability. chang1982diffusion
Si and Ni Optimizing the formation and electrical properties of nickel silicide (NiSi). chang1982diffusion
Si and SiC Investigating Si-SiC interface diffusion to enhance performance in high-power devices. komninou1999gold
Si and Al2O3 Enhancing dielectric properties and adhesion of Al2O3 thin films. werner2011electronic
Si and GaN Examining the impact of interface diffusion on high-frequency devices. lin2000improved
Si and Au Addressing increased contact resistance issues and reliability due to Au diffusion. slezak2000surface
Si and InP Investigating the impact of Si-InP interface diffusion on optoelectronic devices. shapira1984interdiffusion
p- and n-type SCs Exploring methods to minimize ID affecting electrical and thermal conductivities. HeXia
@@
This work was supported by the Science and Technology Development Fund, Macao SAR (File Nos. 0090/2021/A2) and the Guangdong-Hong Kong-Macao Joint Laboratory for Neutron Scattering Science and Technology (Grant No. 2019B121205003).